Solve by substitution:
3x+2y=18
3x+8y=36
Enter answer using parentheses.

Answer by Alan3354(69443)

Solve by substitution:
3x+2y=18
3x+8y=36
-----
3x = -2y + 18
3x = -8y + 36
Sub for 3x
----
-2y + 18 = -8y + 36
6y = 18
y = 3
----
x = 4